# A Review on Deep Learning Autoencoder in the Design of Next-Generation Communication Systems

### v1

@@ Abstract
Traditional mathematical models used in designing next-generation communication systems often fall short due to inherent simplifications, narrow scope, and computational limitations. In recent years, the incorporation of deep learning (DL) methodologies into communication systems has made significant progress in system design and performance optimisation. Autoencoders (AEs) have become essential, enabling end-to-end learning that allows for the combined optimisation of transmitters and receivers. Consequently, AEs offer a data-driven methodology capable of bridging the gap between theoretical models and real-world complexities. The paper presents a comprehensive survey of the application of AEs within communication systems, with a particular focus on their architectures, associated challenges, and future directions. We examine 120 recent studies across wireless, optical, semantic, and quantum communication fields, categorising them according to transceiver design, channel modelling, digital signal processing, and computational complexity. This paper further examines the challenges encountered in the implementation of AEs, including the need for extensive training data, the risk of overfitting, and the requirement for differentiable channel models. Through data-driven approaches, AEs provide robust solutions for end-to-end system optimisation, surpassing traditional mathematical models confined by simplifying assumptions. This paper also summarises the computational complexity associated with AE-based systems by conducting an in-depth analysis employing the metric of floating-point operations per second (FLOPS). This analysis encompasses the evaluation of matrix multiplications, bias additions, and activation functions. This survey aims to establish a roadmap for future research, emphasising the transformative potential of AEs in the formulation of next-generation communication systems.
Index Terms:
Deep learning, Machine learning algorithms, Autoencoder, end-to-end learning, communication systems.
@@ I Introduction
In recent years, the field of communication systems has witnessed significant advancements, particularly with the integration of deep learning techniques. Among these, autoencoders (AEs) have emerged as a powerful tool for designing next-generation communication systems, due to their ability to provide end-to-end optimisation systems, i.e., joint optimisation of both the transmitter and receiver. This paper provides a comprehensive survey on the application of AE in communication systems, focussing on their architectures, challenges, and future directions. The primary objective is to explore how AEs can address the limitations of traditional mathematical models, which often rely on simplifying assumptions and may not capture the full complexity of real-world communication environments. Using the data-driven capabilities of AEs, this study aims to identify the enhancement of the system design and performance of communication systems, offering robust solutions using AEs for end-to-end optimisation.
@@
This review comprehensively investigates the application of AE in communication systems, drawing on 120 recent studies published in reputed publishers, including ScienceDirect, Scopus, Springer and IEEE Xplore, including 2024. The literature search used rigorous keywords such as ”Autoencoder in communication system” and ”deep learning in communication system” to identify relevant studies. By categorising these works based on research issues such as transceiver design, channel models, digital signal processing (DSP) techniques, computational complexity, and non-differentiable channels, we provide a structured overview of the current state of the field. Our analysis encompasses various communication domains, including wireless, optical, semantic, and quantum communication, to identify suitable AE approaches for each. Identified AE approaches are analysed based on their architectures, optimisation methods, layer types, activation functions, and loss functions. In addition, several channel models are investigated, including additive white Gaussian noise (AWGN), multiple-input and multiple-output (MIMO), Rayleigh, and optical channels. Note that many existing studies on AE applications in communication systems rely on simplified channel models such as AWGN or basic fibre models. In addition, this review offers in-depth discussions on existing challenges and potential future research directions. To our knowledge, this is the first comprehensive survey of AE applications in communication systems, providing a valuable roadmap for future research.
@@
This reset of this paper is organised as follows. Section II discusses the challenges of designing next-generation communication systems using traditional models and the advantages of employing AEs. Section III elaborates on the advantages and role of AEs in capturing complex relationships and adapting to real-world scenarios. Section IV outlines the practical challenges of implementing AEs, including the need for extensive training data and the risk of overfitting. Section V describes the end-to-end AE learning process, detailing the architecture and training methods. Section VI explores AE-based communication systems, including symbol-wise and bit-wise approaches, various activation functions in AE models, different loss functions used to optimise AE models, the physical constraints that AEs must consider in communication systems, and different variants of AE architecture and their applications. Section VII delves into geometric and probabilistic constellation shaping techniques. Section VIII focusses on differentiable and non-differentiable channels and components, including how to determine differentiability, proposing solutions for non-differentiable scenarios. Section IX details the calculation of the computational complexity of AE. Section X provides a comprehensive review of the literature on the use of AEs in communication systems, covering wireless, optical, semantic, and quantum communication. Section XI suggests future research areas, and finally Section XII concludes the paper by summarising the potential of AEs to revolutionise communication systems.
II Challenges of Designing Next-Generation Communication Systems
Designing next-generation communication systems based solely on mathematical models may prove insufficient due to several reasons (Fig. 1):
  • Simplifications and assumptions: Mathematical models often rely on simplifying assumptions or approximations to make complex problems more tractable. These simplifications can lead to inaccuracies if the underlying assumptions are not true in practice. For example, Mathematical models for optical fibre design often start with the assumption that light reflects perfectly within the fibre due to total internal reflection, with no losses at the boundaries. Thus, incorporating realistic factors such as scattering, surface roughness, and connector losses into the models is crucial to ensure accurate design and reliable operation [1].
  • Limited scope: Mathematical models may be developed for specific scenarios or conditions by focusing on a specific aspect of the system, such as signal propagation or noise analysis. But they fail to capture the full range of possible operating environments or use cases. For example, when designing optical fibres, ideal properties like uniform refractive index profiles and perfectly circular cores is normally assumed. However, real-world fibres have imperfections due to manufacturing processes, impurities, and microbending losses. These imperfections affect signal propagation, dispersion, and attenuation. Mathematical models may overlook these [2, 3, 4]. On the other hand, as wireless networks increasingly adopt millimeter-wave (mmWave) frequencies, these limitations in scope become more pronounced. Modelling the interaction of signals with environmental factors such as walls, buildings, and even human bodies is crucial for 5G and future systems [5].
  • Lack of physical insight: Over-reliance on mathematical models can lead to a lack of understanding of the underlying physics and mechanisms. This can make it challenging to optimise performance when unexpected behaviour arise [6]. In wireless communication systems, understanding the physical mechanisms behind phenomena such as small-scale fading is critical, particularly in high-mobility scenarios like vehicular networks where doppler shifts play a significant role [7]. In optical communication system, by designing a dense wavelength division multiplexing (DWDM) optical communication system, over-reliance on linear mathematical models can overlook nonlinear effects like four-wave mixing, self-phase modulation, and scattering, leading to unanticipated crosstalk, spectral broadening, and power loss [8]. This lack of physical insight complicates troubleshooting and optimisation, as engineers may struggle to diagnose and mitigate issues like increased BER and reduced transmission distance. Combining models with a deep understanding of the underlying physics is crucial for effective design and maintenance.
  • Inadequate representation of complexity: In wireless communication where a system is operating in environments with high mobility, like long-term evolution (LTE) and 5G networks, it normally can face complex interactions due to handovers, interference, and dynamic spectrum allocation. For example, traditional models may inadequately capture fast-fading environments or the effects of interference in dense deployments of small cells, leading to inaccurate predictions of quality-of-service (QoS) [9]. In optical fibre communication, a system involves complex interactions between light, matter, and energy. Mathematical models may struggle to accurately represent these complexities, leading to inaccuracies in predictions or simulations. For example, mathematical models might inadequately represent the combined impact of chromatic dispersion (CD) and nonlinearities like four-wave mixing, leading to inaccurate predictions of signal degradation over long distances. This simplification can result in underestimating the need for dispersion compensation or additional amplification, ultimately causing performance issues and increased BER in practical deployments [10].
  • Validation challenges: Verifying wireless communication models is particularly challenging in environments with high variability, such as urban or suburban landscapes. For example, testing models for 5G small-cell deployments in different topographies and with varying levels of interference requires extensive field trials or real-world comparisons that are often difficult to scale [11]. In optical fibre communication systems, verifying the accuracy of mathematical models is difficult due to their complexity. Therefore, validation may require extensive experimental testing or comparisons with other models [3, 4, 12].
  • Inadequate consideration of uncertainties: Mathematical models typically assume deterministic behaviour, neglecting the inherent uncertainties and noise present in real-world systems. This can lead to overly optimistic predictions or simulations that do not accurately reflect actual system performance. For example, wireless systems are often subject to signal fading, interference, and fluctuating traffic patterns. Mathematical models for wireless communication may oversimplify by assuming independent, identically distributed channels, whereas in real scenarios, channels are often correlated due to shadowing or mobility [13].
  • Complexity and nonlinearity: Communication systems often exhibit nonlinear behaviour, which can be challenging to model using traditional mathematical techniques due to the difficulty in capturing nonlinear effects. Wireless communication systems, where multi-antenna configurations like massive MIMO are involved, exhibit nonlinear behaviour that can be difficult to model. Interference among multiple users, nonlinearities in radio frequency (RF) components, and beamforming algorithms all introduce complex interactions that challenge traditional linear models [14]. An example in fibre optical communication is a DWDM system with multiple channels where the nonlinear effects (such as self-phase modulation and cross-phase modulation) become significant as the signal power increases. These effects lead to inter-channel crosstalk and impairments that cannot be accurately predicted by simple mathematical models [3, 4, 15]. To account for these nonlinearities, advanced numerical simulations or experimental measurements are necessary .
  • Lack of integration with other disciplines: Mathematical models may not effectively integrate knowledge and insights from related fields such as electrical engineering, materials science, or physics [16]. In wireless communication, it often requires interdisciplinary approaches to combine elements of materials science, such as antenna design, and RF circuit design. Specifically, advances in metamaterials could influence how antennas are modelled, but traditional mathematical approaches might not fully integrate these insights, leading to a gap between theory and practice [17].
MathematicalModels LimitationsLimited scopeSimplificationsand assumptionsLack ofphysical insightInadequate representationof complexityInadequate considerationof uncertaintiesValidation challengesComplexityand nonlinearityLack of integrationwith other disciplines
Figure 1: Communication System Mathematical Model Limitation
Although mathematical models provide valuable information, next-generation communication systems require a holistic approach that considers the above limitations. One promising solution is end-to-end deep learning using AE, which can bridge the gap between mathematical models and real-world complexities, contributing to robust communication systems.
III Why AEs and their Role
An end-to-end learning using AE provides a data-driven approach that complements mathematical models. By capturing intricate relationships and adapting to the complexities of the real world, AEs contribute to the development of robust communication systems. The primary advantages of employing AEs in communication systems include the following.
  • Minimising Distortion: The basic idea behind an AE is to transfer input to output with the least possible distortion, which is perfectly in line with the purpose of a communication system.
  • Framework Alignment: AEs precisely follow the communication system framework, where the encoder acts as the transmitter and the decoder acts as the receiver.
  • Expressive Latent Representation: AEs can create expressive latent representations for complex data, which can be adapted to suit the physical conditions of the communication system.
  • Flexibility: The AE Architectures offer flexibility to be modified for various tasks within communication systems, enhancing their adaptability and efficiency.
Therefore, an end-to-end learning based on AE is an elegant candidate to address the challenges in next-generation communication systems such as:
  • AEs can learn nonlinear mappings between input ( received optical signals) and latent representations. Training an AE on real-world data can capture the complex nonlinearities present in the system. For example, an end-to-end AE can model the nonlinear effects in DWDM systems, such as interchannel crosstalk due to self-phase modulation.
  • AEs can learn compact representations of input data. By encoding and decoding signals, they capture essential features while ignoring noise or imperfections. For example, in fibre design, an AE can learn to represent the refractive index profile of the fibre. It accounts for imperfections (non-uniformity) and helps optimise fibre parameters while considering physical limitations.
  • AEs adapt to changing input distributions. They can learn from historical data and generalise to new conditions. For example, an AE trained on diverse temperature profiles can predict signal behaviour in a temperature-varying fibre link. It can adapt its latent representation to dynamic environments.
IV Challenges of using AE in practical scenarios
Generally the use of AE in practical scenarios can be divided into online and off-line approaches. The off-line approach is that the encoder and decoder parts of the AE model are trained offline based on a real-world dataset. Once training is complete, both can operate independently on the transmitter and receiver sides of the respective communication system. The online approach is that the AE model has partially, or maybe completely, replaced the transmitter and receiver parts of the communication system. This AE model can be pre-trained based on real data-set. There are several challenges when it comes to using AEs in practical scenarios.
  • AE learning from end-to-end needs a significant amount of training data to be effective.
  • AEs can sometimes overfit the training data, which can lead to poor performance with new unseen data. To solve this, a regularisation technique can help prevent overfitting and improve the robustness of the model to noisy data. A powerful technique is to equip each layer with batch normalisation. This normalisation significantly accelerates training and effectively mitigates overfitting [18, 19].
  • AE model may not perform as well in real-world scenarios as it does in the training environment because the AE model may not be able to deal with the unexpected noise or confidence in the input data and practical. There are several methods that can be used to deal with unexpected noisy or corrupted data in practical applications. One approach is to pre-process the data before feeding them into the model, such as cleaning them or removing any obvious outliers. Another option is to use data augmentation techniques, which involve generating multiple variants of the data to increase the amount of training data available to the model. This can help the model learn to handle different types of noise and inconsistencies in the data.
  • An online end-to-end AE learning requires a differentiable channel model to train the transmitter, that is, transmitter training necessitates backpropagating the gradient through the channel. Section VIII-B provides detailed information on potential solutions.
V End-to-end AE Learning
@@
An end-to-end AE is a deep neural network (DNN) model that learns representations (encoding) for sets of data by attempting to reproduce its input in the output. It is an unsupervised machine learning algorithm. This means that it does not require labelled output or targets during training and instead tries to reconstruct inputs from their encoded representation in hidden layers. As shown in Fig. 2, the architecture can be divided into three main parts: Encoder (mapping input $x$ to latent space $w$), Latent Space $w$ (bottleneck), and Decoder (mapping latent space $w$ to output $\hat{x}$).
@@
  1. 1.
    Encoder: The encoder part of the AE takes input data, which can be any form of structured data, and is typically composed of a series of neural network layers with nonlinear activation functions such as rectified linear unit (ReLu) or exponential linear unit (ELU). Its goal is to transform the high-dimensional input into a lower-dimension latent space representation that captures the most important features in an efficient manner.
  2. 2.
    Bottleneck/Latent Space: This part of AE represents compressed knowledge about the original data and is typically composed of one or more un-trainable layers with fewer neurons than the input layer, hence creating a bottleneck effect where only important features are kept while less significant ones get discarded.
  3. 3.
    Decoder: The decoder part of an AE takes the compressed representation and reconstructs the original data from it. It mirrors the encoder architecture but in reverse, attempting to output a reconstruction that is as close as possible to its input (though not necessarily identical due to noise or other factors).
The entire process can be summarised as follows. Input $\mapsto$ Encoding $\mapsto$ Bottleneck/Latent Space $\mapsto$ Decoding $\mapsto$ Output. The AE learns optimal weights during training, which minimises the reconstruction error between inputs and outputs using a loss function. The key advantage of an end-to-end learning approach is that it allows for more efficient data preprocessing or feature extraction since all these tasks are learned simultaneously. It also helps to capture complex patterns within the input data, which can be useful when dealing with high-dimensional datasets.
VI AE based Communication systems
Traditional communication systems employ a block-based approach, where separate transmitter and receiver signal processing blocks are cascaded as shown in Fig. 3 (a). This approach often leads to suboptimal performance due to two main challenges: (1) the difficulty of accurately modelling complex physical channels, as discussed in Section II, and (2) the need to develop specific algorithms to compensate for these effects based on the chosen mathematical model. A promising solution is to use DL-based approaches, inspired by AE architectures [20]. As shown in Fig. 3 (b), by taking advantage of the ability of deep learning models to learn robust latent representations, AE-based systems have the potential to optimise end-to-end communication performance. The latent space representation produced by the encoder is passed through a channel model, where it undergoes channel impairments and produces $y$. Thus, instead of relying on a chain of separate transmitter and receiver signal processing blocks, a data-driven model based on AEs architecture can effectively capture the characteristics of the communication channel, enabling more efficient and reliable transmission.

### v2

@@ Abstract
Traditional mathematical models used in designing communication systems often fall short due to inherent simplifications, narrow scope, and computational limitations. In recent years, the incorporation of deep learning (DL) methodologies into communication systems has made significant progress in system design and performance optimization. Autoencoders (AEs) have become essential, enabling end-to-end learning that allows for the combined optimization of transmitters and receivers. Consequently, AEs offer a data-driven methodology capable of bridging the gap between theoretical models and real-world complexities. This paper presents a comprehensive survey of the application of AEs within communication systems, with a particular focus on their architectures, associated challenges, and future directions. We examine recent studies in the fields of wireless, optical, semantic, and quantum communication, providing a state-of-the-art review with critical analysis. This paper further examines the challenges encountered in the implementation of AEs, including the need for a solution for handling non-differentiable channels, computational complexity, convergence, data representation, and data training, including data scarcity. Moreover, a summary of the computational complexity calculation associated with AE-based systems is provided by conducting an analysis using the metric of floating-point operations per second (FLOPS). Finally, future directions are highlighted to realize the full potential of AEs in creating efficient, reliable, and high-performance communication systems. This survey aims to establish a roadmap for future research, emphasizing the transformative potential of AEs in the formulation of next-generation communication systems.
@@ I Introduction
In recent years, the field of communication systems has witnessed significant advancements, particularly with the integration of deep learning techniques. Among these, autoencoders (AEs) have emerged as a powerful tool for designing next-generation communication systems, due to their ability to provide end-to-end optimization systems, i.e., joint optimization of both the transmitter and receiver. This paper provides a comprehensive survey on the application of AE in communication systems, with a focus on their architectures, challenges, and future directions. The primary objective is to explore how AEs can address the limitations of current communication systems, which often rely on simplifying assumptions and may not capture the full complexity of real-world communication environments.
@@
This review comprehensively investigates the application of AE in communication systems, drawing on recent studies published in reputed publishers, including ScienceDirect, Scopus, Springer, and IEEE Xplore, including 2025. The literature search used rigorous keywords such as ”autoencoder in communication system” and ”end-to-end learning in communication system” to identify relevant studies. By categorizing these works based on research issues such as system design, channel/signal processing, coding/modulation, computational complexity, and non-differentiable channels, we provide a structured overview of the current state of the field. Our analysis encompasses various communication domains, including wireless, optical, semantic, and quantum communication, to identify suitable AE approaches for each. Identified AE approaches are analyzed based on model used, dataset/data type, best performance, and their limitation. Note that many existing studies on AE applications in communication systems rely on simplified channel models such as additive white Gaussian noise (AWGN) or basic fiber models. In addition, this review offers a critical analysis of current work and in-depth discussions on existing challenges and potential future research directions. To our knowledge, this is the first comprehensive survey of AE applications in communication systems, which can provide a valuable roadmap for future research.
@@
This reset of this paper is organized as shown in Fig. 1, where Section II discusses the challenges of designing next-generation communication systems using traditional approaches. Section III elaborates on the advantages of using AEs in communication systems. Section IV describes the end-to-end AE learning process, exploring the AE-based communication architecture, including symbol-wise and bit-wise approaches, different loss functions used to optimize AE models, and the physical constraints that AEs must consider in communication systems. In addition, it delves into geometric and probabilistic constellation shaping techniques and (non)-differentiable channels and components, including how to determine differentiability, proposing solutions for non-differentiable scenarios. Finally, it details the calculation of the computational complexity of AE. Section VI provides a comprehensive review of the literature on the use of AEs in communication systems in various domains, including wireless, optical fiber, free space optical, visible light, semantic, quantum and underwater communications. It also highlights the current challenges for which AE has been used. In addition to the state-of-the-art review, it provides a critical analysis of current approaches highlighting unsolved limitations. Based on that, Section IX suggests future research areas, and finally Section X concludes the paper by summarizing the potential of AEs to revolutionize communication systems.
Section II. Limitations of Mathematical ModelsSimplifications and Narrow ScopeInadequate Handling ofNonlinearity and ComplexityValidation Challenges and Interdisciplinary GapsSection III. Why AEs for Communication Systems?Transmitter/Receiverjoint optimizationPerformance closerto Shannon limitRobustness and adaptabilityto varying channel conditionsFundamental end-to-endcommunication frameworkSection VI. AE Architecture FundamentalsExplaining the simple AE ArchitectureSection V. AE based Communication SystemsAE Architecture approaches:Regression, Symbol-/Bit-wise, Activation function,loss function, physical constraintsAE Architecture Variant:Classical AE, CAE, RAE, VAE, SAE, AAEGeometric and ProbabilisticConstellation ShapingHow to calculateAE ComplexityDomain-Specific Challenges/SolutionsWireless CommunicationsOptical Fiber CommunicationsOptical Wireless CommunicationsSemantic CommunicationsQuantum CommunicationsResearch IssuesSystem DesignChannel/SignalProcessingCoding/ModulationLow ComplexityAE ModelsDense AEVAETurbo AEGAN-AEDenoise AEOther Specific ChallengesComplexity andComputational ConstraintsTraining and ConvergenceData Representation and RateData ScarcityEmerging HW TechnologiesSection VII. (Non-) and Differentiable ChannelHow to determinedifferentiabilitySolutions fornon-differentiableSection VIII. Lessons LearnedRegularization StrategiesLoss Function ChoicesHyperparameter TuningDeployment ReliabilitySection VX. Future Research AreasComplexity andComputational ConstraintsScalability andHigh DimensionsImplementationand HW IssuesNon-differentiable channelsand DSP componentsData and Training IssuesInterpretability and TrustHybrid modelsAdvanced AEarchitecturesSystem-Level andPractical ConsiderationsFederated Learning forDistributed AE TrainingEdge-AI Integration forLow-Latency SystemSection VI. AE-Based Solutions to Domain-Specific Challenges in Communication Systems
Figure 1: The structure of this paper
II Limitations of Mathematical Models
Traditional mathematical models have long served as the basis for designing communication systems. However, their inherent assumptions and simplifications often do not address the complexities of next-generation networks. In the following, we categorize these limitations and highlight how autoencoder (AE)-based approaches offer promising alternatives, albeit with their own challenges.
Simplifications and Narrow Scope
Mathematical models frequently rely on idealized assumptions (e.g., linearity, Gaussian noise) that do not hold in real-world environments. For example, optical fiber models often ignore manufacturing defects or bending losses [1, 2, 3], while wireless channel models oversimplify multipath effects in 5G mmWave systems [4]. AE-based solutions bypass explicit modeling by learning channel characteristics directly from data, which allows for adaptive transceiver optimization (Section V).
Inadequate Handling of Nonlinearity and Complexity
Nonlinear phenomena such as fiber Kerr effects and RF amplifier distortions are poorly captured by analytical models [5, 6, 7]. In wireless systems, understanding effects such as small-scale fading and Doppler shifts is key, especially in high-speed scenarios such as vehicle-to-vehicle networks[8]. Likewise, in dense wavelength division multiplexing (DWDM) optical systems, ignoring nonlinearities such as four-wave mixing and self-phase modulation can result in crosstalk, spectral spreading, and signal degradation[6]. Combining models with physical insight helps engineers better troubleshoot and optimize performance. AE-based solutions excel at modeling nonlinear relationships through deep neural networks (Section IX-3). However, their computational complexity grows significantly with system dimensionality.
Validation Challenges
Testing models in real-world settings, especially for wireless networks in urban or suburban areas, is often resource-intensive. Validating small cell 5G deployments requires extensive trials in varying landscapes and interference levels[9]. Similarly, in fiber optics, the complexity of signal behavior makes it difficult to validate models without experimental setups or detailed benchmarking against known references[2, 3, 10]. Data-driven AE reduce the dependency on closed-form models by training in empirical measurements (Section VI-A1). However, mismatches between training and deployment environments can degrade performance (Section IX-4).
Interdisciplinary Gaps
Mathematical models often operate in isolation, not integrating insights from fields such as materials science or electrical engineering [11]. For example, developments in metamaterials could greatly impact antenna design. But if models do not account for such interdisciplinary advances, they risk lagging behind practical engineering progress [12]. AE-based solution can offer seamless integration with machine learning and signal processing techniques, which allows joint optimization of the physical and protocol layers (Section VI-A2). However, such integration requires careful balance of modularity and end-to-end learning (Section Section IX-6).
III Why AEs for Communication Systems?
Although mathematical models provide valuable information, next-generation communication systems require a holistic approach that considers the above limitations. AEs are increasingly viewed as an elegant solution for next-generation communication systems primarily due to their end-to-end learning capability [13, 14, 15]. This paradigm allows for the joint optimization of the transmitter and receiver based on data. It effectively treats the communication system as an AE aiming to reconstruct the input message at the receiver despite channel impairments. This bypasses the need for explicit, often complex, and idealized mathematical models of the channel and individual processing blocks, which can be particularly challenging in complex or dynamic environments. The end-to-end AE approach has the potential to achieve performance closer to the theoretical Shannon limit by allowing the system to learn optimal joint encoding and decoding strategies directly from the data [16]. The research highlights that end-to-end AE learning can outperform traditional human-engineered transceivers in various scenarios, even without prior knowledge of communication theory [17].
Compared with traditional or hybrid approaches to AE, traditional block-based signal processing, while mature and often achieving near-optimal performance under well-defined conditions, suffers from suboptimal overall system performance due to separate optimization of individual blocks [18]. Traditional methods also rely heavily on explicit probabilistic models that may not adapt well to dynamic channel conditions or capture complex nonlinearities in real world environments [19]. AE-based systems, on the other hand, demonstrate robustness and adaptability to varying channel conditions and can outperform traditional methods, especially in non-linear scenarios where the latter struggle [17, 19, 20, 21]. Furthermore, hybrid approaches have emerged that combine the adaptability of AEs with the efficiency of conventional demapping algorithms to address the drawbacks of purely AE-based systems, such as the high computational complexity and large model sizes associated with training and inference, especially on devices limited by resources [22].
AEs also compare with other machine learning models in their specific applications within communications. Although CNNs are used for tasks such as modulation classification and have shown improvements over traditional expert feature-based methods, AEs offer a more fundamental end-to-end communication framework [13]. GANs are noted for their potential in real-time channel modeling that can be relevant for training end-to-end systems by generating realistic channel impairments, which is relevant for AE training, particularly to address the gradient problem in channel backpropagation [23, 24].
AE-based end-to-end learning is an elegant candidate to address challenges in next-generation communication systems because of its ability to jointly optimize the entire communication link and its data-driven nature, which allows it to handle complex and unknown channel conditions where traditional methods struggle. A compilation of reasons that explain why end-to-end AE-based learning is considered an appropriate candidate includes overcoming the limitations of traditional modular designs [25, 26, 27], handling complex and unknown channel models [28, 29], adaptability and flexibility [30, 31], and simplifying design [13, 26].
IV AE Architecture Fundamentals
@@
An end-to-end AE is a deep neural network (DNN) model that learns representations (encoding) for sets of data by attempting to reproduce its input in the output. It is an unsupervised machine learning algorithm. This means that it does not require labeled output or targets during training and instead tries to reconstruct inputs from their encoded representation in hidden layers. As shown in Fig. 2, the architecture can be divided into three main parts: Encoder (mapping input $x$ to latent space $w$), Latent Space $w$ (bottleneck), and Decoder (mapping latent space $w$ to output $\hat{x}$).
@@
Encoder
The encoder part of the AE takes input data, which can be any form of structured data, and is typically composed of a series of neural network layers with nonlinear activation functions such as rectified linear unit (ReLu) or exponential linear unit (ELU). Its goal is to transform the high-dimensional input into a lower-dimensional latent space representation that captures the most important features in an efficient manner.
Bottleneck/Latent Space
This part of AE represents compressed knowledge about the original data and is typically composed of one or more un-trainable layers with fewer neurons than the input layer, hence creating a bottleneck effect where only important features are kept while less significant ones get discarded.
Decoder
The decoder part of an AE takes the compressed representation and reconstructs the original data. It mirrors the encoder architecture but in reverse, attempting to output a reconstruction that is as close as possible to its input (though not necessarily identical due to noise or other factors).
The entire process can be summarized as follows. Input $\mapsto$ Encoding $\mapsto$ Bottleneck/Latent Space $\mapsto$ Decoding $\mapsto$ Output. The AE learns optimal weights during training, which minimizes the reconstruction error between inputs and outputs using a loss function. The key advantage of an end-to-end AE learning approach is that it allows for more efficient data processing or feature extraction since all these tasks are learned simultaneously. It also helps to capture complex patterns within the input data, which can be useful when dealing with high-dimensional datasets.
V AE-based Communication systems
Traditional communication systems employ a block-based approach, where separate transmitter and receiver signal processing blocks are cascaded as shown in Fig. 3 (a). This approach often leads to suboptimal performance due to two main challenges: (1) the difficulty of accurately modeling complex physical channels, as discussed in Section II, and (2) the need to develop specific algorithms to compensate for these effects based on the chosen mathematical model. A promising solution is to use DL-based approaches, inspired by AE architectures [13]. As shown in Fig. 3 (b), by taking advantage of the ability of deep learning models to learn robust latent representations, AE-based systems have the potential to optimize end-to-end communication performance. The latent space representation produced by the encoder $w$ is passed through a channel model, where it undergoes channel impairments and produces $y$. Thus, instead of relying on a chain of separate transmitter and receiver signal processing blocks, a data-driven model based on AEs architecture can effectively capture the characteristics of the communication channel, enabling more efficient and reliable transmission.